In your English class, essays are 40% of your grade, homework is 20% of your grade, attendance is 5% of your grade, and the final project is 35% of your grade. At the end of the semester, your essay average is 89, your homework average is 90, and your attendance average is 91. What score do you need on the final to get an A- (90%) in the class? (Round your answer to 2 decimal places.)
Answer #1

Solution. English class, essays are 40% of your grade, homework is 20% of your grade, attendance is 5% of your grade, and the final project is 35% of your grade. At the end of the semester, your essay average is 89, your homework average is 90, and your attendance average is 91. What score do you need on the final to get an A- (90%

= 88(40/100)+91(20/100)+95(5/100)+final(35/100)

= 3520+1820+475 + final 35=9000

final project= 3185/35= 91

Thank you